OpenStudy (anonymous):

factor the expression 7x^2-46x-21

OpenStudy (anonymous):

Use magic x

OpenStudy (anonymous):

or maybe quadratic formula...

OpenStudy (anonymous):

Do you have trouble starting the problem?

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

use quadratic formula

OpenStudy (anonymous):

7x^2 - 46x - 21(1) Use the factoring AC Method,( You Tube). Find 2 numbers b1 and b2 that satisfy 2 conditions: b1* b2 = a*c = -147, and (b1 + b2) = -46. Compose factor of a*c = 147: (1, -147)(3, -49). Then, b1 = 3 and b2 = -49. Replace in the trinomial (1) the term -46x by the 2 terms (3x) and (-49x), then factor.

OpenStudy (anonymous):

Correction: compose factor pairs of a*c = -147: (1, -147)(3, -49)

OpenStudy (anonymous):

7x^2 - 46x - 21= 0 (1). Factoring is the same as solving quadratic equation. I show you a new method, the Transforming Method (on Google Search) that solves very fast this type of equation. It proceeds through 3 steps. Step 1. Transform this equation into a simplified form x^2 + bx + a*c = 0 (2) with a = 1 and with C = a*c. Step 2. Solve equation (2) by the Diagonal Sum method that immediately obtain the 2 real roots. Roots have different signs (Rule of signs). Compose factor pairs of a*c = -147. Proceeding: (-1, 147)(-3, 49). This last sum is -3 + 49 = 46 = -b. Then, the 2 real roots of (2) are y1 = -3 and y2 = 49. Step 3. Back to the original equation (1), the 2 real roots are: x1 = y1/7 = -3/7, and x2 = y2/7 = 49/7 = 7. Factoring form; (x - 7)(7x + 3)